After many tests, we now think either the dome encoder is failing or the dome power supply is incorrectly set, following the powerout for the UPS replacement. More updates soon.
The pointing was recentered (off after worm-gear lubrication work) and the focus was readjusted (for seasonal drift).
Several images are blank, likely with the dome completely in the way
The centering of the field seems to drift from night to night
At high hour angle, there is significant trailing in the images. There is still notable “coma” in images near zenith
Several have noticed the following behavior in the last few nights with the RCT:
I believe these are all symptoms of the encoder slippage and worm gear oscillation. I believe they needing cleaning/re-greasing. I’ve put in a service request with KPNO. I still think part of the PSF is mirror support related, but we wont know for sure until we eliminate the other sources of confusion.
Tonight I was able to zero point the telescope, and test the guider. It appears that the pointing tolerance for accurate guiding is about 15″, or ~20 pix from the absolute center of the chip. Guiding works better now, but it does not improve the PSF substantially.
The camera is hung in a state that requires a power cycling.
Camera fails to execute a software reset
The camera returns the error “failure in reset”
The leach controller window shows highly incorrect values in gain (e.g., -4095), amp, heat, etc.
Cycle the power on the camera on site (cannot be done remotely). It is unclear where this problem stems from.
Cloud limit was raised to 70%. This should allow some backlogged requests to get through.
There was an unknown error with the leach controller daemon last night. The daemon was restarted, and now appears to be working.
Please ignore the nomets messages in last night’s queue. We were running a test. Weather was to poor to observe.
The chip orientation has been erroneous since moving to amp=C a long while ago. The chip orientation should now be correct with North=Left East=Down.
Several images the last few nights are blank. I suspect the dome encoders are off, and have re-homed the encoder. Please send comments if your images are still blank.
There is a bug in the RCT scripts that occurs when parsing names with white spaces from the JPL Horizons ephemeris service. I’m still working on robust solutions. You may try putting quotes around your object names, like “C/2011 L1”. But I recommend looking up targets on http://ssd.jpl.nasa.gov/horizons.cgi to generate a Object Data Page, and using either the Record Number (indicated by Rec #: xxxxx in the upper left hand corner) or the MB number (in the upper right hand corner) as the object name in the IsPlanet key on the RCT Observing Scheduler.
Telescope was adjusted for seasonal drift in focus. Images now show a three-point PSF indicicating that the mirror support compressor is not opperating properly.